No.1 Cab overhaul is progressing steadily. Much of the wiring is
complete now and requires final testing. The push buttons on the desk
have been replaced wholesale with examples extracted from scrap Class
56s which were in exemplary condition compared to those in D1516, with
little dirt and much better contact faces. After a light clean and
check, the "new" switches have been attached behind the bezels which
themselves have been fully dismantled and stripped prior to repainting
in matt grey. The wires which the switches control have been
re-terminated and the cable (the only original as-built cable in the
locomotive) has been checked for insulation integrity. Much time has
been absorbed in the re-painting of the desk and panels, which have now
been fully finished with a coat of matt varnish to preserve the finish
(we hope!). The instrument bezels have been refinished in satin black.
The steam gauge bezel will require re-working to accommodate it in the
desk as will the screenwash cover at the side of the desk. The reason
for this being that these items do not belong to 1516, and were sourced
to replace items lost during her period of storage before moving to the MR-B. The
air tight (allegedly!) seals between all panels have been replaced with
new self adhesive foam which meets full RGS (Railway Group Standards)
standards for fire etc. this is being used in place of the original
felt which was removed to repaint the base metal of the desk. Also the
rubber kick strip at the base of the removable cab panels is being
replaced where the original material has worn through.
rest of the locomotive is being worked on simultaneously. Much of the
team's attention is now focused on the generator bay of the engine
room, cleaning back the old paint and applying anti-corrosive primer
prior to it being glossed. Some re-panelling will be necessary in the
floor area to repair the thinner areas of the loco's floor which have fallen
victim to forty odd years of exposure. This activity is in anticipation
of the overhauled generator's return which will allow the continuation
of the stripping of the power unit. Thus far the locking nuts and
gears have been removed from the majority of big ends, the big ends
themselves following when we can bar the power unit over again (which
requires the generator).
Meanwhile a full complement of
overhauled cylinder heads awaits refitting to the locomotive. These are
presently in storage where they continue to demolish our shins as we
try to stagger past them in the dark!
The auxiliary machines are
being prepared for dispatch to Messrs. Dowding and Mills for attention
to their respective electrical machines. Thus far the compressors have
been mechanically disconnected and require de-wiring to complete the
job. Photos of the ongoing work on the cab desk can be seen on our
47417 Gallery Page.
47401 SEPTEMBER 2005 UPDATE
The final gala of 2005 presented a scenario that the team had started to become strangers to. The former D1500 was seeing out the year in her inimitable manner and wouldn’t start. This
was a bit odd since most of the normal suspects that could prevent her
from starting had been steadily rectified through our ongoing program
to increase electrical reliability, and ultimately the ERV (Engine Run
Valve) was energised so the last line of defence/irritation in the
power units electrical control equipment was operating. The
ERV, which sits in the governor assembly is energised when the various
protective interlocks in the loco are satisfied that the power unit can
be safely started. When ERV is energised it allows oil
pressurised by the “triple pump” to open the fuel pump racks via a
hydraulic actuator, in much the same manner as you would depress the
throttle of a car to start it. Thus all things being equal it should have started – if there was sufficient oil pressure. Out of the top five of Brush 4 failure modes, an expired triple pump nylon gearwheel ranks high. Sure enough the oil pump drive gearwheel resembled a pulley. The pump was removed and a spare gear (removed from the pump in stores) grafted onto the known good pump. This
was necessary as the pump in store was found to have an erratic action,
which felt like a damaged bearing race and could have resulted in the
spare pump gearwheel being stripped of its teeth instantly. The
Thursday evening before the gala saw the pre-filled pump fitted and
whipped up – and predictably she didn’t start – what a surprise.
Bleeding of the oil system upwards from the triple pump was carried out on Friday. When
the start button was depressed the racks opened wide, as they should
before pressure diminished and the governor returned the racks to the
rest position. This still looked like an air lock (but looked curable) hence the comment on the website “she may run” - sorry we were so vague! Come the Saturday we were quietly confident (read “excessively optimistic”). There were two feasible approaches. Firstly try the union at the back of the governor for one last bleed. Secondly
it was realised that the batteries may need a big dose of charge as the
pump would run OK until the engine started cranking and the battery
volts (and therefore amps) would drop significantly, reducing the speed
of the electrically driven triple pump and allowing the racks to close
down again as oil pressure faltered.
of you around on Saturday will know the answer – one and a half hours
of charge, one last bleed up and we hit the button – and it worked. She took power at both ends without fault and all with half an hour to spare before being due off shed – result!
JUNE 2005 UPDATE - D1516's CYLINDER HEADS BEFORE AND AFTER!
Three of the overhauled cylinder heads have been returned to Butterley. The photos (dated 24th June 2005) can be seen here showing
the second batch of heads at Fragonset, Derby, prior to work being
carried out and on the right are the three completed heads. Pistons and
cylinder liners were due to be the next items to be removed from
has begun on rewiring the No. 1 end cab desk where a number of cables
to one of the switch panels had been severed, refurbishing the switch
panel itself and replacing the wiring to the cab foot warmers.
47401 PROGRESS REPORT - JULY 2005
most recent running day was on Sunday 19th June, the loco completing
five round trips without incident. A further run was to have taken
place on Saturday 2nd July on a vice steam turn but the loco had to be
stood down owing to no trained driver being available.
work has included attention to a leaking gauge on the header tank. We
have also replaced the STR relay after it failed several weeks ago
meaning that the engine can once again be shut down without resort to
the fuel rack.
D1516 CYLINDER HEADS REMOVED 7th MAY 2005
restoration of D1516 took a further step forward with the removal of
the loco's cylinder heads which will be sent away to Derby for
overhaul. The photos, here, show
one of the heads being craned off the power unit plus a collection of
six of the heads (half way there!) on the ground, and also the power
unit following removal of the heads and awaiting removal of pistons and
47401 OIL CHANGE & RUNNING DAY 10th APRIL 2005
oil change was carried out on 47401 over the weekend of 2nd/3rd April,
the first time this has been carried out in preservation. Approx. 150
gallons of waste oil had to be drained out into 5 gallon drums while
the loco stood over the pit at Swanwick, then £400 worth of new oil was
slowly pumped in - this should see us through for another few years
operation. Thanks to everyone who turned out this weekend to assist!
was running on 10th April at the MRB's Diesel Gala. A minor fault
kept the loco on shed on the Saturday, 9th April, however the booked
diagram was successfully completed on the Sunday without major incident!
MARCH 2005 - D1516 UPDATE
the previous report on 47417 the interiror of the engine room has been
cleaned and painted prior to receiving the overhauled generator.
Elsewhere on the loco the battery isolating switch has been reinstalled
in the boiler room and the surrounding area repainted. A selection of
photographs taken on 12th March 2005 showing the work done and some of
the jobs still to be tackled including severe corrosion in the roof and
silencer section as well as cab desk rewiring can be found here (scroll down to the photos dated 12th March 2005).
SUTTON-IN-ASHFIELD - FEBRUARY 2005
visit to inspect the work carried out on D1516's generator at the
premises of Messrs Dowding & Mills, Sutton in Ashfield, was made
during February 2005. Photos can be seen on the 47417 Gallery Page